Perform an appropriate hypothesis test to determine if there is a significant relationship between a student’s...

  1. Perform an appropriate hypothesis test to determine if there is a significant relationship between a student’s race and the punishment received for talking back to teachers in a school district. Interpret the findings.

In School Suspension

Out of School Suspension

Intermediate Punishment

African-American/Black

19

52

9

Hispanic/Latino

14

39

7

White

9

12

24
Answer #1

Following table shows the row total and column total:

In school suspension Out of school suspension Intermediate punishment Total
African-American/Black 19 52 9 80
Hispanic/Latino 14 39 7 60
White 9 12 24 45
Total 42 103 40 185

Expected frequencies will be calculated as follows:

Following table shows the expected frequencies:

In school suspension Out of school suspension Intermediate punishment Total
African-American/Black 18.162 44.541 17.297 80
Hispanic/Latino 13.622 33.405 12.973 60
White 10.216 25.054 9.73 45
Total 42 103 40 185

Following table shows the calculations for chi square test statistics:

O E (O-E)^2/E
19 18.162 0.038665565
14 13.622 0.010489209
9 10.216 0.144739233
52 44.541 1.249111627
39 33.405 0.937105972
12 25.054 6.801585216
9 17.297 3.979892987
7 12.973 2.750075464
24 9.73 20.9283556
Total 36.84002088

Following is the test statistics:

Degree of freedom: df =( number of rows -1)*(number of columns-1) = (3-1)*(3-1)=4

The p-value is: 0.0000

Conclusion: Since p-value is less than 0.05 so we reject the null hypothesis. That is we can conclude that there is a significant relationship between a student’s race and the punishment received for talking back to teachers in a school district.


Excel function used for p-value: "=CHIDIST(36.84, 4)"